OpenStudy (anonymous):

solve and check your answer. 4/1-x=5/1+x

OpenStudy (anonymous):

hello this equation has this structure a / b = c / d ----------> a (d) = b (c)

OpenStudy (anonymous):

ok ?

OpenStudy (mathlegend):

4/1 and the 5/1 supposed to be fractions?

OpenStudy (anonymous):

yes

OpenStudy (mathlegend):

We know that any number over 1 is just itself... so lets rewrite this okay? 4-x=5+x

OpenStudy (mathlegend):

I like to keep the variable positive. 4-x=5+x +x +x ----------- 4=5+2x

OpenStudy (mathlegend):

(Now, lets work on getting the variable on one side and the regular numbers on the other side of the equal sign) 4 =5+2x -5 -5 -------- -1=2x x=-1/2

OpenStudy (anonymous):

ok thanks :)